# Predicting Subway Passenger Flows under Incident Situation with Causality

### v1

@@ 3.2.1 Feature selection
It is important to note that certain indicators, such as the delay time caused by an incident, the number of trains delayed by five minutes or more and the overall duration of the incident, can only be obtained after the incidents have concluded. Consequently, these indicators cannot be accessed when predicting causal effects. However, we believe this is not a significant issue, as experienced staff in the subway system can generally extrapolate these indicators based on the type or other characteristics of the incident. Additionally, the severity of the incident must be considered when predicting its causal effects; therefore, we have no choice but to include these indicators. In future work, features related to the types of incidents and the degree of damage to facilities could be considered as indicators of incident severity, and these features can be obtained immediately after an incident occurs. A description of the selected features is provided in Table 1.
@@ 3.2.2 Sample selection
According to the formal section, we can calculate the causal effects for all OD pairs and all time intervals for an incident. However, these differences are not solely attributable to the incident; some may be influenced by random effects. For example, an incident situation is show in Figure 3, and this incident may clearly impact the passenger flow from station A to station B, but the effect on the passenger flow from station C to station B may be less apparent. Although the difference between the counterfactual and observed passenger flow from station C to station B can indeed be calculated, it is essential to determine whether this difference is due to random effects or the incident itself. Therefore, it is necessary to select the significantly affected passenger flows, as illustrated in the upper part of Figure 4.
@@
$y=\epsilon_{1}$ with probability $1-p$;
@@
$E(\beta-\hat{\beta})^{2}=(1-p)^{2}\ |\beta|^{2}+c\frac{p\sigma_{2}^{2}+\sigma_{1}^{2}}{n}$
@@
It can be founded that the distance is bound by two terms, the first term is $\left(1-p\right)\left|\beta\right|^{2}$, which means the larger the $p$ is, the smaller the loss will be. This law is very intuitive because that if more samples in the dataset are actually affected by incidents, the fitting results will be better. The last term $\frac{p\sigma_{2}^{2}+\sigma_{1}^{2}}{n}$ is related to sample size $n$, which means that the more samples are in the dataset, the better the fitting result will be. This law is also intuitive. In summary, the theorem means that we need to consider both the sample quantity ($n$) and the sample quality ($p$) when selecting which sample should be used to fit the prediction model. As a result, if we just use all OD records after incidents, then the sample quantity $n$ could be large but sample quality $p$ might be small. Therefore, if we can propose a method to judge which causal effect is more likely to be true effect and which is more likely to be randomness, then we can select these true effects and control the sample quality $p$ to be larger and obtain a better fitting result.
@@ 3.2.3 Significant test
This null hypothesis means the actual OD flow under incident $x_{i,k}^{1}$ is the same as the counterfactuals OD flow in normal situation $x_{i,k}^{0}$. Therefore, the calculated causal effect, i.e., $x_{i,k}^{1}-{\hat{x}}_{i,k}^{0}$, is because of randomness. We use a placebo test to construct the test statistic [2] and determine whether to reject the null hypothesis. If the null hypothesis is rejected, then we can regard this causal effect is different from zero significantly. The placebo test is widely used, distribution-free and similar to the permutation test. We will elaborate it as follows.
[…]
@@ 3.3 Prediction adjustment
We use the following strategy to adjust OD flow prediction: If there are many ODs and these ODs are affected by an incident with different probability according to relationship between these ODs and the incident. Then we note the probability of a specific OD to be affected by the incident is $p$. Then the strategy is: First determining a threshold $P$ (which is $21-p$ 2 in Figure 4), for each OD, if the probability of the OD to be affected by the incident, i.e., $p$, is greater than $P$, we will use causal effect prediction model to adjust it, otherwise we will just regard the prediction of normal model as the final prediction value, as shown in Figure 4. Then we can obtain the risk of this strategy.
@@
Where $\epsilon_{1}$ following normal distribution $N(0,\sigma_{1}^{2})$ and $\epsilon_{2}$ following normal distribution $N(0,\sigma_{2}^{2})$, $\beta$ is the parameter and $p$ is a constant between 0 and 1. If there are a lot of samples $\left(x_{i},p_{i}\right)$ and a prediction model
$\hat{y}=\hat{f}(x)$
Then we can get the risk of the above strategy.
$E\left(y-\hat{y}\right)^{2}=\frac{1}{2}P^{2}(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2})-PE\hat{f}\left(x\right)^{2}+c$
@@
Besides, the probability of a specific OD to be affected by the incident, i.e., $p_{i}$, is not known, but we can use the p-value in history data (calculated by significant test section) and some influence factors to fit a model to predict $p$. As a result, the workflow of training and deploying the causal effect prediction model could be shown in Figure 4, and $p1,p2$ in the figure are thresholds used to select training samples and samples need correction.
@@ 4.1 Data description
Most incidents in our dataset had minimal impact, resulting in no or insignificant delays. Consequently, we only selected those incidents that caused delays exceeding five minutes. Additionally, we excluded incidents that occurred during the early morning hours, such as those involving the first train of the day, as passenger flow during these times is inherently low, rendering the analysis less meaningful. Ultimately, we identified six eligible incidents, with the last incident in our dataset occurring on March 14th. Our objective is to utilize the data prior to March 14th to build models that predict passenger flow on that day. It is noted that the incident began at 7:15 a.m. and ended at 9:24 a.m.; therefore, our focus is to predict passenger flows between 7:00 a.m. and 12:30 a.m. Given that the incident occurred in suburban sections, OD pairs whose origin and destination stations are not located on the incident line are unlikely to be affected. Therefore, our analysis concentrates on ODs with at least one station (either the origin or destination) located on the incident line.
@@ 4.2 Causal effect identification results
The effect of an incident typically increases with its onset and gradually dissipates as the incident ends. Consequently, we analyze all OD pairs from the beginning of an incident until three hours after its end. Using the model proposed in the Method section, we identify the causal effects of each subway incident and conduct significance tests on these effects. However, due to space limitations, we only provide a detailed report on the causal effect identification results for the incident in 11th Marth. The details of this incident are as follows:
Starting at 7:45 on the 11th Marth, a vehicle malfunction occurred on Line 11, and the malfunctioning vehicle was arranged to proceed slowly. From 7:51 to 8:43, temporary passenger flow control measures were implemented at Malu, Chenxiang Highway, Nanxiang, Taopu Xincun, Qilianshan Road, Liziyuan, and Fengqiao Road stations.
@@
Figure 5 illustrates the significantly affected ODs identified during the six time periods from 8:00 AM to 10:30 AM. In the figure, the blue line represents the subway line where the incident occurred, the black lines denote regular subway lines, and the red lines indicate the ODs that were significantly impacted by the incident.
It should be noted that the incident occurred at 7:45 AM and ended at approximately 8:45 AM. As illustrated in Figure 5, the number of affected ODs initially increased and peaked between 8:30 and 9:00 AM, before gradually declining. By 10:00 and 10:30 AM, only a few ODs remained affected. Additionally, at the onset of the incident, relatively few ODs were impacted, as the effects had not yet spread widely.
Furthermore, the affected stations are primarily located in the northern part of Shanghai. Figure LABEL:fig confirms that the ODs affected by this incident are predominantly situated in the northern region, which is consistent with the actual situation.
@@
First, it is important to note that March 11th is a Monday, and the following Monday is March 18th. Additionally, there is no data for March 16th due to missing data in the original dataset. Each subplot in Figure 7 illustrates the changing trend of an OD, with the gray box on March 11th indicating the time of the incident.
For these four ODs, the patterns observed during the incident period differ significantly from those on other days. For example, the top subplot, which illustrates the OD from Jiading West to Chenxiang Road, indicates that the OD value within the gray box is considerably higher compared to other weekdays, such as March 12th, March 13th, and the following Monday, March 18th. Similar patterns can be observed in the other three ODs. This phenomenon may be attributed to the primary consequence of the incident: delays occurring around 8:00 AM. These delays could lead passengers who would typically take the subway at 8:00 AM to board later, while those who planned to travel after 8:00 AM continued with their original schedules. This overlap of passengers resulted in a heightened peak during the morning rush hour.
@@ 4.3 Passenger flow prediction results
We selected two matrix-based models: ST-ResNet [50] and UNet [31], which are widely utilized for matrix data, and GEML [41], a model specifically designed for Ride-hailing OD prediction.
We also conducted experiments with different causal effect prediction models. Specifically, we selected three models: Linear Regression, Random Forest, and Gradient Boosted Decision Trees. Following [54], we omitted the ODs with passenger flow less than 2. Besides, MAE (mean absolute err), RMSE (root mean square error) and MAPE (mean absolute percentage error) are used to evaluate the prediction accuracy. The equations to calculate them are in follows.
$MAE=\frac{1}{n}\sum{|\hat{y}-y|}$
$RMSE=\sqrt{\frac{\sum{\hat{y}-y)^{2}}}{n}}$
$MAPE=\frac{1}{n}\sum{\frac{{|\hat{y}-y}|}{y}}$
@@
First, we summarize the performance of these baseline models without adjustments for causal effects. The errors on all passenger flows and significantly affected passenger flows (p-values $\leq$ 0.05) show that most models meet a substantial increase in error when predicting the flows of significantly affected ODs. In these cases, errors even surge by as much as three times. Traditional machine learning methods generally produce higher prediction errors across all flows compared to deep learning models. However, for significantly affected flows, traditional models exhibit lower prediction errors than deep learning models. This discrepancy may arise because deep learning models capture subtle and sensitive data patterns that may not be applicable under incident conditions, resulting in rapid increases in error when deployed in such scenarios.
@@ 4.4 Analysis of causal effect prediction model
Table 3 shows that the counterfactual passenger flow ($x_{0}$) is the most significant element affecting the causal effect of incidents, which is consistent with common sense. This demonstrates that if an OD has a higher baseline passenger flow, an incident will have a greater influence on that flow. In contrast, for OD with smaller baseline flows, the incident exhibits a modest effect it. Other key factors include the time from the start and end of the incident to the time of OD, which also corresponds to common expectations. If the interval is too short, there may not be sufficient time for the incident to influence passenger flow; if it is too long, the impact might have already dissipated. Therefore, the time interval is crucial to the effect of the incident. Additionally, the delay duration caused by the incident is also highly important, which makes sense, as longer delays generally reflect more severe incidents, and therefore, a greater impact on passenger flow. However, features such as the number of canceled trains and the number of five-minute delay trains appear less significant. This might be because the severity of incidents does not directly correspond to these metrics. For instance, the count of five-minute delay trains might increase as the incident becomes more serious, but as the severity intensifies further, some delayed trains might be canceled, causing the number to drop. The lack of a one-to-one relationship suggests that this metric may not clearly reflect the severity of incidents. These findings indicate that future models may benefit from including indicators that reflect the incident severity more accurately.
@@
Moreover, we analyzed model outputs by plotting partial dependence plots (PDPs) using the SHAP method [33]. Each subplot in Figure 8 represents a PDP and shows the interaction between two variables, where the x-axis indicates one variable and the color represents another. For example, in the first subplot, larger counterfactual baseline flows ($x_{0}$) generally correlate with stronger causal effects. However, the signs of the causal effect vary by color, which reflects the time difference between the OD and the start of the incident. Smaller differences (bluer points) tend to have negative causal effects, indicating suppressed flow, while larger differences (redder points) suggest increased flow, as some users delay travel and lead to a recovery effect.
@@
The first and second subfigures in the second row illustrate how the distance from the incident location to the origin and destination stations affects the causal effects. The effect is generally stronger when the distance is within approximately 10, but it diminishes significantly beyond this range. When the distance is less than 10, the absolute values of the causal effect are relatively consistent, suggesting a threshold effect of distance. Additionally, the distance to the origin station has a more pronounced influence on the causal effect than the distance to the destination station.

### v2

@@ 3.2.1 Feature selection
It is important to note that certain indicators, such as the delay time caused by an incident, the number of trains delayed by five minutes or more and the overall duration of the incident, can only be obtained after the incidents have concluded. Consequently, these indicators cannot be accessed when predicting causal effects. However, we believe this is not a significant issue, as experienced staff in the subway system can generally extrapolate these indicators based on the type or other characteristics of the incident. Additionally, the severity of the incident must be considered when predicting its causal effects; therefore, we have no choice but to include these indicators. In future work, features related to the types of incidents and the degree of damage to facilities could be considered as indicators of incident severity, and these features are avaliable immediately after an incident occurs. A description of the selected features is provided in Table 1.
@@ 3.2.2 Sample selection
According to the previous section, we can calculate the causal effects for all OD pairs and all time intervals for an incident. However, these differences are not solely attributable to the incident; some may be influenced by random effects. For example, an incident situation is show in Figure 3, and this incident may clearly impact the passenger flow from station A to station B, but the effect on the passenger flow from station C to station B may be less apparent. Although the difference between the counterfactual and observed passenger flow from station C to station B can indeed be calculated, it is essential to determine whether this difference is due to random effects or the incident itself. Therefore, it is necessary to select the significantly affected passenger flows, as illustrated in the upper part of Figure 4.
@@
If the influence factors about the causal effects is noted as $x$, the causal effect is denated as $y$, assume $y$ and $x$ obeys the following rule:
$y=\epsilon_{1}$ with probability 1p1-p;
@@
Where $\epsilon_{1}$ follows normal distribution $N(0,\sigma_{1}^{2})$ and $\epsilon_{2}$ follows normal distribution $N(0,\sigma_{2}^{2})$, $\beta$ is the parameter and $p$ is a constant between 0 and 1.
@@
$E(\beta-\hat{\beta})^{2}=(1-p)^{2}\ |\beta|^{2}+c\frac{p\sigma_{2}^{2}+\sigma_{1}^{2}}{n}$ (6)
@@
It can be found that the distance is bound by two terms, the first term is $\left(1-p\right)\left|\beta\right|^{2}$, which means the larger the $p$ is, the smaller the loss will be. This law is consistent with normal intuition because if more samples in the dataset are actually affected by incidents, the fitting results will be better. The last term $\frac{p\sigma_{2}^{2}+\sigma_{1}^{2}}{n}$ is related to sample size $n$, which means that the more samples are in the dataset, the better the fitting result will be. This law is also intuitive. In summary, the theorem means that we need to consider both the sample quantity ($n$) and the sample quality ($p$) when selecting which sample should be used to fit the prediction model. As a result, if we just use all OD records after incidents, then the sample quantity $n$ could be large but sample quality $p$ might be small. Therefore, if we can propose a method to judge which causal effect is more likely to be true effect and which is more likely to be randomness, then we can select these true effects and control the sample quality $p$ to be larger and obtain a better fitting result.
@@ 3.2.3 Significant test
This null hypothesis means the actual OD flow under incident $x_{i,k}^{1}$ is the same as the counterfactuals OD flow in normal situation $x_{i,k}^{0}$. Therefore, the calculated causal effect, i.e., $x_{i,k}^{1}-{\hat{x}}_{i,k}^{0}$, is because of randomness. We use a placebo test to construct the test statistic [Abadie2010SyntheticCM] and determine whether to reject the null hypothesis. If the null hypothesis is rejected, then we can regard this causal effect is different from zero significantly. The placebo test is widely used, distribution-free and similar to the permutation test. We will elaborate it as follows.
Recall the method from the causal effect estimation section, we use $(a_{i,k},x_{i,k})$ for $i\in\ \{1,2,\ldots,l-1,l+1,\ldots,n\}$ and $a_{l,k}$ to infer $x_{l,k}^{0}$. Let $f$ represent this process. Besides, let $S$ denote the set $\{(a_{i,k},\ x_{i,k})\mid\ i\in\{1,\ 2,\ldots,n\}\}$, and $S_{-i}$ denote the set obtained by removing $(a_{i,k},\ x_{i,k})$ from $S$. Thus, we have:
@@ 3.3 Prediction adjustment
We use the following strategy to adjust OD flow prediction: If there are many ODs and these ODs are affected by an incident with different probability according to relationship between these ODs and the incident. Then we note the probability of a specific OD to be affected by the incident is $p$. Then the strategy is: First determining a threshold $P$ (which is 1p21-p2 in Figure 4), for each OD, if the probability of the OD to be affected by the incident, i.e., $p$, is greater than $P$, we will use causal effect prediction model to adjust it, otherwise we will just regard the prediction of normal model as the final prediction value, as shown in Figure 4. Then we can obtain the risk of this strategy.
@@
Where $\epsilon_{1}$ follows normal distribution $N(0,\sigma_{1}^{2})$ and $\epsilon_{2}$ follows normal distribution $N(0,\sigma_{2}^{2})$, $\beta$ is the parameter and $p$ is a constant between 0 and 1. If there are a lot of samples $\left(x_{i},p_{i}\right)$ and a prediction model $\hat{y}=\hat{f}(x)$. Then we can get the risk of the above strategy.
$E\left(y-\hat{y}\right)^{2}=\frac{1}{2}P^{2}(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2})-PE\hat{f}\left(x\right)^{2}+c$ (10)
@@
Moreover, the probability that a specific OD pair is affected by an incident ($p_{i}$) is not directly observable. However, we can utilize historically derived p-values (calculated in the significance test section) along with relevant influence factors to develop a predictive model for $p_{i}$. In our actual experimental implementation, we employ a random forest model and use identical predictors in the causal effect prediction model to predict the significance level of causal effects. As a result, the workflow of training and deploying the causal effect prediction model could be shown in Figure 4, and $p1,p2$ in the figure are thresholds used to select training samples and samples need correction.
@@ 4.1 Data description
Besides, it should be emphasized that exit-time OD cannot be used to generate bridge bus directly. In practice, exit-time OD needs to be subtracted from OD travel time to transform into entry-time OD. Moreover, if entry-time OD data become accessible in the future, developing models for entry-time OD should be attempted.
@@
Most incidents in our dataset had minimal impact, resulting in no or insignificant delays. Consequently, we only selected those incidents that caused delays exceeding five minutes. Additionally, we excluded incidents that occurred during the early morning hours, such as those involving the first train of the day, as passenger flow during these times is inherently low, rendering the analysis less meaningful. Ultimately, we identified six eligible incidents, with the last incident in our dataset occurring on March 14th. Our objective is to utilize the data prior to March 14th to build models that predict passenger flow on that day. It is noted that the incident began at 7:15 a.m. and ended at 9:24 a.m.; therefore, our focus is to predict passenger flows between 7:00 a.m. and 12:30 p.m. Given that the incident occurred in suburban sections, OD pairs whose origin and destination stations are not located on the incident line are unlikely to be affected. Therefore, our analysis concentrates on ODs with at least one station (either the origin or destination) located on the incident line.
@@ 4.2 Causal effect identification results
The effect of an incident typically increases with its onset and gradually dissipates as the incident ends. Consequently, we analyze all OD pairs from the beginning of an incident until three hours after its end. Using the model proposed in the Method section, we identify the causal effects of each subway incident and conduct significance tests on these effects. However, due to space limitations, we only provide a detailed report on the causal effect identification results for the incident in Marth 11th. The details of this incident are as follows:
Starting at 7:45 on the Marth 11th, a vehicle malfunction occurred on Line 11, and the malfunctioning vehicle was arranged to proceed slowly. From 7:51 to 8:43, temporary passenger flow control measures were implemented at Malu, Chenxiang Highway, Nanxiang, Taopu Xincun, Qilianshan Road, Liziyuan, and Fengqiao Road stations.
@@
Figure 5 illustrates the significantly affected ODs identified during the six time periods from 8:00 a.m. to 10:30 a.m.. In the figure, the blue line represents the subway line where the incident occurred, the black lines denote regular subway lines, and the red lines indicate the ODs that were significantly impacted by the incident.
It should be noted that the incident occurred at 7:45 a.m. and ended at approximately 8:45 a.m.. As illustrated in Figure 5, the number of affected ODs initially increased and peaked between 8:30 and 9:00 a.m., before gradually declining. By 10:00 and 10:30 a.m., only a few ODs remained affected. Additionally, at the onset of the incident, relatively few ODs were impacted, as the effects had not yet spread widely.
Furthermore, the affected stations are primarily located in the northern part of Shanghai. Figure 5 confirms that the ODs affected by this incident are predominantly situated in the northern region, which is consistent with the actual situation.
@@
First, it is important to note that March 11th is a Monday, and the following Monday is March 18th. Additionally, there is no data for March 16th due to missing data in the original dataset. Each subplot in Figure 7 illustrates the changing trend of an OD, with the red box on March 11th indicating the time of the incident.
For these four ODs, the patterns observed during the incident period differ significantly from those on other days. For example, the top subplot, which illustrates the OD from Jiading West to Chenxiang Road, indicates that the OD value within the red box is considerably higher compared to other weekdays, such as March 12th, March 13th, and the following Monday, March 18th. Similar patterns can be observed in the other three ODs. This phenomenon may be attributed to the primary consequence of the incident: delays occurring around 8:00 a.m.. These delays could lead passengers who would typically take the subway at 8:00 a.m. to board later, while those who planned to travel after 8:00 a.m. continued with their original schedules. This overlap of passengers resulted in a heightened peak during the morning rush hour.
@@ 4.3 Passenger flow prediction results
We employed 5-fold cross-validated grid search to determine optimal parameters for RF and GBDT models. For the RF implementation, we evaluated n_estimators across [100, 300, 500], max_depth across [5, 10, 15] and min_samples_split across [2, 5, 10]. For the GBDT framework, we optimized n_estimators within [200, 500, 800], learning_rate across [0.01, 0.05, 0.1], max_depth in [3, 5, 7], and subsample ratio between [0.8,0.9, 1.0].
We selected two matrix-based models: ST-ResNet [Zhang_Zheng_Qi_2017] and UNet [ronneberger2015u], which are widely utilized for matrix data, and GEML [Wang2019OriginDestinationMP], a model specifically designed for Ride-hailing OD prediction.
We also conducted experiments with different causal effect prediction models. Specifically, we selected three models: Linear Regression, Random Forest, and Gradient Boosted Decision Trees. Following [Zou2024RealtimePO], we omitted the ODs with passenger flow less than 2. Besides, MAE (mean absolute err), RMSE (root mean square error) and MAPE (mean absolute percentage error) are used to evaluate the prediction accuracy. These metrics are calculated as follows:
$MAE=\frac{1}{n}\sum{|\hat{y}-y|}$ (12)
$RMSE=\sqrt{\frac{\sum{(\hat{y}-y)^{2}}}{n}}$ (13)
$MAPE=\frac{1}{n}\sum{\frac{{|\hat{y}-y}|}{y}}$ (14)
@@
First, we summarize the performance of these baseline models without adjustments for causal effects. The errors on all passenger flows and significantly affected passenger flows (p-values $\leq$ 0.05) show that most models experience a substantial increase in error when predicting the flows of significantly affected ODs. In these cases, errors even surge by as much as three times. Traditional machine learning methods generally produce higher prediction errors across all flows compared to deep learning models. However, for significantly affected flows, traditional models exhibit lower prediction errors than deep learning models. This discrepancy likely occurs because deep learning models capture subtle and sensitive data patterns. However, these patterns often fail to generalize during incident conditions. Consequently, when deployed in such scenarios, the models exhibit rapid error increases.
@@ 4.4 Analysis of causal effect prediction model
Table 3 shows that the counterfactual passenger flow ($x_{0}$) is the most significant element affecting the causal effect of incidents, which is consistent with common sense. This demonstrates that if an OD has a higher baseline passenger flow, an incident will have a greater influence on it. In contrast, for OD with smaller baseline flow, the incident exhibits a modest effect on it. Other key factors include the time from the start and end of the incident to the time of OD, which also corresponds to common expectations. Short intervals leave insufficient time for impacts to manifest, while long intervals allow effects to dissipate. Therefore, the time interval is crucial to the effect of the incident. Additionally, the delay duration caused by the incident is also highly important, which makes sense, as longer delays generally reflect more severe incidents, and therefore, a greater impact on passenger flow.
However, features such as the number of canceled trains and the number of five-minute delay trains appear less significant. This might be because the severity of incidents does not directly correspond to these metrics. For instance, the count of five-minute delay trains might increase as the incident becomes more serious, but as the severity intensifies further, some delayed trains might be canceled, causing the number to decrease. The lack of a one-to-one relationship suggests that this metric may not clearly reflect the severity of incidents. These findings indicate that future models may benefit from including indicators that reflect the incident severity more accurately.
@@
Besides, we conducted ablation experiments to validate the importance of each feature group on causal effect prediction. The results are represented in Tab 4. The ablation study reveals that removing any feature group degrades model accuracy; the most significant performance degradation occurs when excluding counterfactual passenger flow features. This aligns with our feature importance analysis, confirming that counterfactual passenger flow ($x_{0}$) is the most critical predictor of incident impacts.
R2 MAE RMSE
Ori model 0.550 4.909 7.605
[…]
Table 4: Ablation experiment results
Moreover, we analyzed model outputs by plotting partial dependence plots (PDPs) using the SHAP method [scott2017unified]. Each subplot in Figure 8 represents a PDP and shows the interaction between two variables, where the x-axis indicates one variable and the color represents another. For example, in the first subplot, larger counterfactual baseline flows ($x_{0}$) generally correlate with stronger causal effects. However, the signs of the causal effect vary by color, which reflects the time difference between the OD and the start of the incident. Smaller time differences (bluer points) correlate with negative causal effects, indicating suppressed flow, while larger time differences (redder points) suggest increased flow, as some users delay travel and lead to a recovery effect.
@@
The first and second subfigures in the second row illustrate how the distance from the incident location to the origin and destination stations affects the causal effects. The effect is generally stronger when the distance is within approximately 10, but it diminishes significantly beyond this range. When the distance is less than 10, the absolute values of the causal effect are generally small, suggesting a threshold effect of distance. Additionally, the distance to the origin station has a more pronounced influence on the causal effect than the distance to the destination station.